How many ways can 7 beads be placed on a bracelet with no claps?

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

7 ways to choose the first*6*5*...*1=5040 ways

Answer by ikleyn(52780)

.
How many ways can 7 beads be placed on a bracelet with no claps?
~~~~~~~~~~~~~~~~~~~~~~~~~~~


I assume that it is the same  (or similar)  as to place  7  persons around a circular table.

The answer is  (7-1)! = 6! = 1*2*3*4*5*6 = 720.

For  7  items,  there are  6!  distinguishable circular permutations.

--------------------

To see many other similar  (and different)  solved problems,  look into the lesson
    - Persons sitting around a circular table
in this site,  and learn the subject from there.


Precisely  THIS  given problem was solved there,  TOO.  See problem  6  of that lesson.